OpenStudy (silverlark2):

a force of 10N is applied on a 2kg metal block for a time of 3s. What is the rate of change of momentum?

OpenStudy (anonymous):

10 Newtons.

OpenStudy (anonymous):

F= ma = mv/t a = 10/2 = 5 mv/t = 2*5 = 10 N